OpenStudy (anonymous):

Plz Help!! Find the slope of a line on a graph (-2,-1), (-4,4)

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

i know u gave me the formula i dont understand it though

OpenStudy (anonymous):

you can totally do this! \[\frac{ y-y_1 }{ x-x_1 }\] just plug in the numbers...

OpenStudy (anonymous):

um....i cant lol im not smart

OpenStudy (anonymous):

(-2,-1) and (-4, 4)...let's start with what are your y-coordinates?

OpenStudy (anonymous):

coordinates/ordered pairs are always written as (x,y).

OpenStudy (anonymous):

1 and 4

OpenStudy (anonymous):

check again...negatives matter

OpenStudy (anonymous):

-1,4?

OpenStudy (anonymous):

bingo! ok, so let's plug that in...\[\frac{ y-y_1 }{ x-x_1 }=\frac{ -1-4 }{ x-x_1 }\]

OpenStudy (anonymous):

now let's figure out our x-coordinates...

OpenStudy (anonymous):

then -2,4

OpenStudy (anonymous):

your coordinates are (-2,-1) and (-4,4) so I agree with -2 being an x-coordinate, but what's the first number in the 2nd set of coordinates?

OpenStudy (anonymous):

@Whalexnuker

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

ok, so this is the tricky part...

OpenStudy (anonymous):

\[\frac{ -1-4 }{ -2--4 }\]

OpenStudy (anonymous):

do you see how -1 is the first number on top and -2 is the first number on the bottom? That's because those 2 numbers go together (-2,-1)

OpenStudy (anonymous):

there is only one -4

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

the denominator (bottom) says -2 minus -4

OpenStudy (anonymous):

oh lol srry

OpenStudy (anonymous):

it's ok :) do you know how to subtract negative numbers?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

after this i have one more

OpenStudy (anonymous):

-2-(-4) isn't going to equal -6. that would be true if you were adding, but you're subtracting.

OpenStudy (anonymous):

subtracting a negative is like adding a positive...so instead of -2-(-4) think -2+4

OpenStudy (anonymous):

oh so then 2

OpenStudy (anonymous):

perfect, what about the top number?

OpenStudy (anonymous):

-1-4

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

so put it all together and you have your slope! -5/2

OpenStudy (anonymous):

thnk you so much

OpenStudy (anonymous):

you are so very welcome! I wish I could stay and help more, but I have work in the morning. Good luck on the rest of your problems!